OpenStudy (anonymous):

Pararllel to y=1/2x+9 going through the point -4,2

OpenStudy (anonymous):

2= .5(-4) + b 2= -2+b add 2 to both sides 4 = b then y=1/2x+4 Do you get how to do it now?

OpenStudy (anonymous):

Since the two lines are parallel, their slopes are same i.e. 1/2 ( Equation of line is of the form y=mx+c, where c is the y intercept) Now plug in the values of x and y you have got in the equation y=(1/2)*x+c and find the value of c which is 4. Now you have the equation of the line as y=(1/2)*x+4

OpenStudy (jdoe0001):

hmmm what's the slope of y=1/2x+9 ?